# Next Generation Equation-Free Multiscale Modelling of Crowd Dynamics via Machine Learning

### v2

@@ 4 Numerical results
The comparison between model classes reveals that the use of MVARs provides a more accurate and reliable framework for long-term predictions than that of LSTMs. This is clearly demonstrated by their significantly lower errors and narrower error percentiles in the closed-loop setting (Table 3, Fig. 5). Crucially, this performance gap widens substantially when compared to the open-loop results (Table 2, Fig. 4), where LSTMs performed more competitively. For instance, the MVAR(9) model’s mean $L_{2}$ error across time increased from $\sim$7% (open-loop) to $\sim$14% (closed-loop) within the “baseline” reconstruction error from POD and perturbed initial conditions, respectively, while the LSTM(9) model’s error changing respectively from $\sim$19% to $\sim$35% is well above these “baselines”. This superior performance of MVARs is attributed to the challenge of error accumulation in closed-loop forecasting; the linear structure of the MVAR model proves to be much less susceptible to this instability than complex, nonlinear LSTMs. While LSTMs can theoretically capture richer nonlinear dynamics, their performance severely degrades due to the “curse of dimensionality” in their training, i.e., finding a global optimum, accompanied by a distribution shift [99, 100] (i.e., the error accumulation of the recursive scheme during inference, that shifts the distribution upon which the models were trained), despite their ability to learn the one-step dynamics effectively. In addition, there are several challenges regarding the “correct” initialization of LSTMs [101], especially in dynamical systems and multiscale modeling tasks. In contrast, the training of the linear MVAR models is based only on OLS, thus giving a global optimum. For the particular applications, the results suggest that MVARs are a much better choice for long-term horizon predictions in the latent space. This is a result in line also with other studies (see e.g., [62]).
Beyond accuracy, the proposed framework offers a balanced trade-off between prediction accuracy and computational efficiency for coarse-grained crowd dynamics. In particular, for the benchmark case of $100$ pedestrians moving in the corridor for $250\text{s}$, the conventional numerical integration requires a computational time of $20\text{s}$ for the microscopic SFM simulator simulation plus $30\text{s}$ for the extraction of the density fields (in a coarser time discretization), summing up to a total of $50\text{s}$. In contrast, our framework requires just $0063\text{s}, 0.0$$0161\text{s}$ for the MVARs models and $6.75\text{s}$, $6.77\text{s}$ for the LSTM ones to complete the predictions across the long-time horizon integration–representing a speed-up of 3,100x and 7.4x for the MVARs and LSTMs, respectively. Even when accounting for the offline costs (POD basis construction and ROM training), MVARs retain a significant advantage, with training times under $0.005\text{s}$ (as shown in Table 1). The MVARs variants prove particularly efficient for real-time applications, while maintaining better accuracy than their LSTM counterparts. These gains are especially valuable for real-world scenarios, where the number of pedestrians renders the conventional approaches infeasible.
@@ 5 Conclusions
In this work, we present a four-step data-driven framework that combines manifold and machine learning to construct a discrete evolution operator for emergent crowd dynamics. Our framework provides a computationally efficient and systematic approach for learning ROMs for spatio-temporal emergent crowd dynamics. By first encoding the full high‐dimensional state into a compact latent representation, we learn the system’s evolution operator using surrogate ROMs, bypassing the prohibitive computational cost of learning PDEs via, for example, CNNs or other high‐dimensional ML models such as DNNs. Working in latent space with delayed POD coordinates simplifies the enforcement of physical constraints (such as mass conservation) over long time horizons, since, as we demonstrate, the POD algorithm used to find an appropriate set of coordinates of the latent space explicitly conserves the mass of the reconstructed dynamics. In fact, under the usual generic assumptions of Takens’ theorem, applying time-delay embeddings to POD modes produces a delay-coordinate latent model that is diffeomorphic to the original system dynamics (see also results and discussion in [67, 62, 102, 68, 69]).
@@
Thus, our approach does not attempt to learn any closed‑form PDE, but instead learns the effective solution operator of such an unavailable PDE, mapping initial crowd density to its future evolution. The resulting ROMs trained in the latent space, in the form of linear (MVARs) or nonlinear multivariate autoregressive models (such as LSTMs), encode and reconstruct this operator numerically via the POD reconstruction. As a result, our framework allows for rapid and accurate “what‐if” scenarios exploration and real‐time forecasting of crowd dynamics. Critically, we demonstrate that linear MVAR models within our framework not only outperform more complex LSTMs in long-term predictive accuracy and reliability, but also achieve a computational speed-up of over three orders of magnitude compared to conventional simulation. This combination of interpretability, speed, and accuracy makes our approach, particularly with MVARs, an efficient approach for large‐scale, real-world crowd‐dynamics analysis and distributed control [103, 104, 19, 21, 24, 22, 105].
Finally, our finding that simple linear MVAR models outperform nonlinear LSTM networks in the latent space in terms of complexity, training time, and interpretability is a significant insight. It aligns with a growing body of literature [62, 102, 69, 63] that demonstrates the power of linear ROMs constructed in judiciously chosen latent spaces (e.g., POD-based delay-coordinate embeddings, Koopman operator) for approximating complex nonlinear or chaotic dynamics, offering a promising path toward interpretable and efficient data-driven modeling.
@@ Appendix A Social Force Model
where $C_{i}\in\mathbb{R}^{+}$ is the strength of the repulsive interaction against a wall or an obstacle, $D_{i}$ is its corresponding decay parameter, and $d_{iB}$ is the euclidean distance between the pedestrian and the closest point of the object $B$; $\bm{n}_{iB}$ is the respective the unit vector. Unlike pedestrian–pedestrian interactions, no tangential (sliding) friction is typically included for static obstacles, as they do not exerft dynamic resistance.
@@
Appendix B Initial Conditions of Microscopic Distributions
    $p(x_{i},y_{i})=\mathcal{U}\left([x_{\text{min}},x_{\text{max}}]\times[y_{\text{min}},y_{\text{max}}]\right).$ (B.1)
    $p(x_{i},y_{i})=\frac{1}{2\pi\sigma_{x}\sigma_{y}}\exp\left(-\frac{(x_{i}-\mu_{x})^{2}}{2\sigma_{x}^{2}}-\frac{(y_{i}-\mu_{y})^{2}}{2\sigma_{y}^{2}}\right),$ (B.2)
    $p(x_{i},y_{i})$ $=\frac{1}{\sqrt{2\pi}\,\sigma_{y}}\,\exp\!\left(-\frac{(y-\mu_{y})^{2}}{2\sigma_{y}^{2}}\right)$ (B.3)
    $p(x_{i},y_{i})=\frac{1}{\pi s_{x}s_{y}}\cos\left(\frac{x_{i}-c_{x}}{s_{x}}\right)\cos\left(\frac{y_{i}-c_{y}}{s_{y}}\right).$ (B.4)
The training and testing data sets were generated from $C=20$ initial conditions, with 5 samples drawn from each of the four distributions described above. To ensure plausible and varied crowd configurations, the parameters for these distributions were selected heuristically to provide a broad spatial coverage while mitigating boundary artifacts. Furthermore, during the sampling process, pedestrian positions were resampled if they were within a critical distance equal to $2r_{i}$ of an existing pedestrian or a wall/obstacle to prevent initial overlaps and collisions. The specific parameter configurations used for the training and testing data sets are detailed Table B.1 and Table B.2, respectively.
Table B.1: Microscopic initial distribution configurations and parameters for constructing the training dataset.
Case c Distribution Type Parameters
1 Uniform, Eq. (B.1) $x_{\text{min}}=2,x_{\text{max}}=15,y_{\text{min}}=3,y_{\text{max}}=9$
2 Uniform, Eq. (B.1) $x_{\text{min}}=33,x_{\text{max}}=46,y_{\text{min}}=3,y_{\text{max}}=9$
3 Uniform, Eq. (B.1) $x_{\text{min}}=2,x_{\text{max}}=20,y_{\text{min}}=4,y_{\text{max}}=8$
4 Uniform, Eq. (B.1) $x_{\text{min}}=28,x_{\text{max}}=46,y_{\text{max}}=4,y_{\text{max}}=8$
5 Uniform, Eq. (B.1) $x_{\text{min}}=3,x_{\text{max}}=17,y_{\text{max}}=4,y_{\text{max}}=8$
6 Gaussian, Eq. (B.2) $\mu_{x}=10,\;\mu_{y}=6,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
7 Gaussian, Eq. (B.2) $\mu_{x}=10,\;\mu_{y}=6,\sigma_{x}=2,\;\sigma_{y}=1.5$
8 Gaussian, Eq. (B.2) $\mu_{x}=38,\;\mu_{y}=6,\sigma_{x}=2,\;\sigma_{y}=1.5$
9 Gaussian, Eq. (B.2) $\mu_{x}=38,\;\mu_{y}=6,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
10 Gaussian, Eq. (B.2) $\mu_{x}=37,\;\mu_{y}=5,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
11 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=8,\;\mu_{y}=7,\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=12,\;\sigma_{x_{2}}=0.8$
12 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=10,\;\mu_{y}=6,\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=14,\;\sigma_{x_{2}}=1$
13 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=28,\;\mu_{y}=6,\sigma_{x_{1}}=1.3,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=32,\;\sigma_{x_{2}}=1.1$
14 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=36,\;\mu_{y}=6,\sigma_{x_{1}}=1.4,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=40,\;\sigma_{x_{2}}=1.2$
15 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=9,\;\mu_{y}=6,\sigma_{x_{1}}=0.8,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=11,\;\sigma_{x_{2}}=1$
16 Cosine, Eq. (B.4) $c_{x}=10,\;c_{y}=6,s_{x}=10,\;s_{y}=3$
17 Cosine, Eq. (B.4) $c_{x}=20,\;c_{y}=8,s_{x}=10,\;s_{y}=3$
18 Cosine, Eq. (B.4) $c_{x}=30,\;c_{y}=7,s_{x}=10,\;s_{y}=3$
19 Cosine, Eq. (B.4) $c_{x}=40,\;c_{y}=6,s_{x}=10,\;s_{y}=3$
20 Cosine, Eq. (B.4) $c_{x}=10,\;c_{y}=6,s_{x}=9,\;s_{y}=3$
Table B.2: Microscopic initial distribution configurations and parameters for constructing the testing dataset.
Case c Distribution Type Parameters
1 Uniform, Eq. (B.1) $x_{\text{min}}=2,\;x_{\text{max}}=16,\;y_{\text{min}}=3,\;y_{\text{max}}=10$
2 Uniform, Eq. (B.1) $x_{\text{min}}=32,\;x_{\text{max}}=46,\;y_{\text{min}}=3,\;y_{\text{max}}=10$
3 Uniform, Eq. (B.1) $x_{\text{min}}=2,\;x_{\text{max}}=19,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
4 Uniform, Eq. (B.1) $x_{\text{min}}=29,\;x_{\text{max}}=46,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
5 Uniform, Eq. (B.1) $x_{\text{min}}=4,\;x_{\text{max}}=17,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
6 Gaussian, Eq. (B.2) $\mu_{x}=12,\;\mu_{y}=7,\;\sigma_{x}=2,\;\sigma_{y}=1.5$
7 Gaussian, Eq. (B.2) $\mu_{x}=11,\;\mu_{y}=5,\;\sigma_{x}=1.5,\;\sigma_{y}=2$
8 Gaussian, Eq. (B.2) $\mu_{x}=38,\;\mu_{y}=6,\;\sigma_{x}=2,\;\sigma_{y}=1.5$
9 Gaussian, Eq. (B.2) $\mu_{x}=39,\;\mu_{y}=5,\;\sigma_{x}=1.5,\;\sigma_{y}=1.5$
10 Gaussian, Eq. (B.2) $\mu_{x}=37,\;\mu_{y}=7,\;\sigma_{x}=1.5,\;\sigma_{y}=1.5$
11 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=7,\;\mu_{y}=7,\;\sigma_{x_{1}}=1.1,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=11,\;\sigma_{x_{2}}=0.9$
12 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=10,\;\mu_{y}=7,\;\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=14,\;\sigma_{x_{2}}=1$
13 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=29,\;\mu_{y}=6,\;\sigma_{x_{1}}=1.3,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=33,\;\sigma_{x_{2}}=1.1$
14 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=37,\;\mu_{y}=6,\;\sigma_{x_{1}}=1.4,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=41,\;\sigma_{x_{2}}=1.2$
15 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=10,\;\mu_{y}=6,\;\sigma_{x_{1}}=0.8,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=12,\;\sigma_{x_{2}}=1$
16 Cosine, Eq. (B.4) $c_{x}=15,\;c_{y}=4,\;s_{x}=8,\;s_{y}=3$
17 Cosine, Eq. (B.4) $c_{x}=25,\;c_{y}=6,\;s_{x}=10,\;s_{y}=3$
18 Cosine, Eq. (B.4) $c_{x}=35,\;c_{y}=5,\;s_{x}=10,\;s_{y}=3$
19 Cosine, Eq. (B.4) $c_{x}=45,\;c_{y}=7,\;s_{x}=5,\;s_{y}=3$
20 Cosine, Eq. (B.4) $c_{x}=15,\;c_{y}=6,\;s_{x}=7,\;s_{y}=3$
Appendix C Density fields extraction via Kernel Density Estimation
@@
$K_{h}(\mathbf{x}-\mathbf{x}_{i}(t))=\frac{1}{2\pi\det(\mathbf{H})^{1/2}}\exp\left(-\tfrac{1}{2}(\mathbf{x}-\mathbf{x}_{i}(t))^{\top}\mathbf{H}^{-1}(\mathbf{x}-\mathbf{x}_{i}(t))\right),$ (C.1)
where $\mathbf{H}$ is the symmetric positive-definite bandwidth matrix, $\mathbf{x}\in\mathbb{R}^{2}$ is the spatial domain and $\mathbf{x}_{i}(t)\in\mathbb{R}^{2}$ is the position of the $i$-th pedestrian.
@@
$h_{i}=\sigma_{i}\left(\frac{4}{N(d+2)}\right)^{1/(d+4)},$ (C.2)
@@
$\Omega^{\text{ext}}=[a-\Delta x,b+\Delta x]\times[c,d],$ (C.3)
while the support in the $y$-direction remains unbounded. To enforce continuity and reduce boundary-induced artifacts in the $x$-direction, we further augment the observed positions with “ghost” particles near the domain edges. Let $\epsilon=\Delta x$ define a threshold for proximity to the boundary. For each position $\mathbf{x}_{i}(t)=(x_{i}(t),y_{i}(t))$ such that $x_{i}(t)\leq a+\epsilon$, we introduce a “mirrored” pedestrians at
$\mathbf{x}_{i}^{\text{aug1}}(t)=(x_{i}(t)-(b-a),y_{i}(t)).$
Similarly, for every $\mathbf{x}_{i}(t)$ such that $x_{i}(t)\geq b-\epsilon$, we introduce a “mirrored” pedestrians at
$\mathbf{x}_{i}^{\text{aug2}}(t)=(x_{i}(t)+(b-a),y_{i}(t)).$
@@
$\{\mathbf{x}_{i}^{\prime}(t)\}=\{\mathbf{x}_{i}(t)\}\cup\{\mathbf{x}_{i}^{\text{aug1}}(t)\}\cup\{\mathbf{x}_{i}^{\text{aug2}}(t)\}.$ (C.4)
A pseudo-code for the estimation of the density in $\Omega\subset\mathbb{R}^{2}$ is given in Algorithm C.1.
Algorithm C.1 Kernel Density Estimation (KDE) for extracting density fields from pedestrian positions.
@@
4:Define extended support $\Omega^{\text{ext}}$ to handle periodic boundary conditions $\triangleright$ Use Eq. (C.3)
@@
6:Introduce “mirrored” pedestrians to form $\{\mathbf{x}_{i}^{\prime}(t)\}$ $\triangleright$ Use Eq. (C.4)
@@
9:  Compute kernel contribution $K_{h}(\mathbf{x}-\mathbf{x}_{p}(t))$ over the extended support $\Omega^{\text{ext}}$ $\triangleright$ Use Eq. (C.1)
@@
14:Employ binary mask function to ensure zero density in obstacle subdomain $\Gamma$, by setting $\triangleright$ See Eq. (C.5)
@@
$M(x,y)=\begin{cases}0,&\text{if }(x,y)\in\Gamma\subset\Omega,\\ 1,&\text{otherwise}.\end{cases}$ (C.5)

### v3

@@ 4.2 Counterflow case study
The comparison between model classes in both unidirectional and counterflow case studies reveals that the use of MVARs provides a slightly more accurate and reliable framework for long-term predictions than that of LSTMs. This is clearly demonstrated by their lower errors and narrower error percentiles in the closed-loop setting (Table 2, Fig. 4 and Table F.2, Fig. 5). While all models perform remarkably well in an open-loop setting (see Table E.1, Fig. E.1 and Table F.2, Fig. F.2), the effect of error accumulation in recursive predictions degrades their predictive accuracy. For instance, in the unidirectional flow case the MVAR(4) model’s mean $L_{2}$ error across time increased from $\sim$4% (open-loop) to $\sim$15% (closed-loop), and the LSTM(4) model’s error increased from $\sim$4% to $\sim$16%. This consistent pattern in both case studies with different latent dimensions ($6$ vs $24$) and interaction complexities is attributed to the challenge of error accumulation in closed-loop forecasting; the linear structure of the MVAR model proves to be less susceptible to this instability than complex, nonlinear LSTMs. While LSTMs can theoretically capture richer nonlinear dynamics, their performance degrades due to the “curse of dimensionality” in their training, i.e., finding a global optimum, accompanied by a distribution shift [20, 54] (i.e., the error accumulation of the recursive scheme during inference, that shifts the distribution upon which the models were trained), despite their ability to learn the one-step dynamics effectively. In addition, there are several challenges regarding the “correct” initialization of LSTMs [70], especially in dynamical systems and multiscale modeling tasks. In contrast, the training of the linear MVAR models is based on least squares regression, thus giving a global optimum. For the particular applications, the results suggest that MVARs are a better choice for long-term horizon predictions in the latent space. This is a result in line also with other studies (see e.g., [91]).
The proposed framework offers a balanced trade-off between prediction accuracy and computational efficiency for coarse-grained crowd dynamics.
For the unidirectional flow case ($250\,\text{s}$ horizon, $K=1000$ steps), one conventional simulation requires $80\,\text{s}$ for SFM integration plus approximately $18\,\text{s}$ for density field extraction via KDE, totaling $98\,\text{s}$ per complete prediction. Our framework’s online closed-loop prediction requires $w\times(t_{ID}+t_{R})\,\text{s}$ for density extraction and restriction of the first $w$ lagged snapshots, followed by $(1000-w)\times(t_{F}+t_{L})$ for per-step ROM forecasting and lifting operations; see Eq. (51). This results in total online execution times of $0.397$, $0.887$, $1.413$, and $1.871\,\text{s}$ for the MVAR(4), MVAR(9), LSTM(4), and LSTM(9) models, respectively, corresponding to speed-ups of $247\times$, $110\times$, $69\times$, and $52\times$ relative to the conventional approach. For the counterflow case, one conventional simulation requires a total of $77.35\,\text{s}$, while our framework yields total online execution times of $0.781$ and $2.141\,\text{s}$ for the MVAR(10) and LSTM(10) models, respectively, corresponding to speed-ups of $99\times$ and $36\times$. These online times include lifting at every time step for a fair comparison with the conventional simulation output. In practice, with on-demand lifting (i.e., applying $\mathcal{L}$ only when density fields are needed), the speed-ups would be even greater, as the framework evolves in the latent space (Eq. (51)).
Offline costs (one-time) for the unidirectional flow case include $(80+18)\times 20=1960\,\text{s}$ for generating $20$ training trajectories via SFM with KDE extraction, $4.24\,\text{s}$ for POD basis computation, and model training times of $0.0013$, $0.0017$, $62$, and $86\,\text{s}$ for MVAR(4), MVAR(9), LSTM(4), and LSTM(9) models respectively (the LSTMs times are averaged over 50 runs; see Table 1). For the counterflow case, the offline costs are $1547\,\text{s}$ for generating $20$ training trajectories, $15.51\,\text{s}$ for the augmented POD basis, and model training times of $0.031$ and $104\,\text{s}$ for MVAR(10) and LSTM(10) models respectively (see Table F.1). While offline costs appear significant, they become negligible when amortized over multiple predictions. For example, the total offline cost for LSTM(9) in the unidirectional flow case ($2050.24\,\text{s}$) equals just $21$ conventional simulations; a cost quickly recouped when dozens or hundreds of predictions are required, as in parameter studies or real-time applications.
Overall, MVAR variants prove particularly efficient for real-time applications, while maintaining better accuracy than their LSTM counterparts. These gains are especially valuable for real-world scenarios, where the number of pedestrians renders the conventional approaches infeasible.
@@ 5 Conclusions
In this work, we present a four-step data-driven framework that combines manifold and machine learning to construct a discrete evolution operator for emergent crowd dynamics. Our framework provides a computationally efficient and systematic approach for learning ROMs for spatio-temporal emergent crowd dynamics. By first encoding the full high‐dimensional state into a compact latent representation, we learn the system’s evolution operator using surrogate ROMs, bypassing the prohibitive computational cost of learning PDEs via, for example, CNNs or other high‐dimensional ML models such as DNNs. Working in latent spaces with delayed POD coordinates simplifies the enforcement of physical constraints (such as mass conservation) over long time horizons, since, as we demonstrate, the POD algorithm used to find an appropriate set of coordinates of the latent space explicitly conserves the mass of the reconstructed dynamics. In fact, under the usual generic assumptions of Takens’/ Whitneys’ embeddings theorems, applying time-delay embeddings to POD modes produces a delay-coordinate latent model that is diffeomorphic to the original system dynamics (see also results and discussion in [91, 102, 71, 9, 44]). For our illustrations, we assessed our approach using data generated from the SFM in a corridor with an obstacle under both unidirectional and counterflow scenarios, though more advanced crowd models incorporating behavioral aspects or optimal path planning [15, 34] can be readily employed.
@@
Our approach does not attempt to learn any closed‑form PDE, but instead learns the effective solution operator of such an unavailable PDE, mapping initial crowd density to its future evolution. The resulting ROMs trained in the latent space, in the form of linear (MVARs) or nonlinear multivariate autoregressive models (such as LSTMs), encode and reconstruct this operator numerically via the POD reconstruction. As a result, our framework allows for rapid and accurate “what‐if” scenarios exploration and real‐time forecasting of crowd dynamics. Critically, we demonstrate that linear MVAR models within our framework not only outperform more complex LSTMs in long-term predictive accuracy and reliability, but also achieve a computational speed-up of over two orders of magnitude compared to conventional simulation. This combination of interpretability, speed, and accuracy makes our approach, particularly with MVARs, an efficient approach for large‐scale, real-world crowd‐dynamics analysis and distributed control [36, 5, 8, 53, 27, 6, 82]. The predictive performance of linear MVAR models in our framework (consistently matching or exceeding that of nonlinear LSTMs) aligns with a growing body of literature [91, 49, 9, 44] demonstrating that linear ROMs constructed in judiciously chosen latent spaces (e.g., POD-based delay-coordinate embeddings, Koopman operator) can effectively approximate complex nonlinear or chaotic dynamics. For example, nonlinear manifold learning techniques such as Diffusion Maps have been shown to provide more accurate embeddings of complex dynamics compared to POD, particularly during transients [37].
As formulated in its vanilla form, the proposed method inherits the well-known limitations of POD-based reduced-order models, and in particular their strong dependence on the geometries and boundary conditions represented in the training data, and the present algorithm is no exception: if the test configuration departs substantially away from the snapshot ensemble, the linear subspace may fail to represent the solution manifold accurately. Handling boundary-aware cases would require a different modeling approach, such as NOs. Incorporating NOs in latent spaces within the proposed framework is a promising direction which will be investigated in future studies, including a broader comparison with alternative surrogate models as for example via Transformer-based architectures [60, 101] that have shown a good performance for long-horizon modeling, and latent Neural ODEs [98], which offer a continuous-time perspective for learning reduced dynamics. Moreover, approaches based on Learning Effective Dynamics (LED) [71, 24, 109, 42] and its extensions, such as G-LED [50] and interpretable LED [84], provide complementary frameworks that emphasize structure preservation and interpretability of the learned operators. A vital extension of this work for real-world deployment is the transition to probabilistic forecasting to quantify predictive uncertainty. While techniques such as Monte Carlo dropout or quantile regression could be applied, the linear Gaussian structure of the MVAR model offers a particularly direct and interpretable pathway to achieve this by propagating initial condition uncertainty through its stochastic latent dynamics. Finally, the proposed mass-conserving framework provides a foundation for applications to real-world trajectory data, systematic benchmarking against alternative surrogates, and extension to other collective motion systems such as vehicular traffic.
@@ Appendix A Social Force Model
where $C_{i}\in\mathbb{R}^{+}$ is the strength of the repulsive interaction against a wall or an obstacle, $D_{i}$ is its corresponding decay parameter, and $d_{iB}$ is the euclidean distance between the pedestrian and the closest point of the object $B$; $\bm{n}_{iB}$ is the respective the unit vector. Unlike pedestrian–pedestrian interactions, no tangential (sliding) friction is typically included for static obstacles, as they do not exert dynamic resistance.
@@
Appendix B Density fields extraction via Kernel Density Estimation
@@
$K_{h}(\mathbf{x}-\mathbf{x}_{i}(t))=\frac{1}{2\pi\det(\mathbf{H})^{1/2}}\exp\left(-\tfrac{1}{2}(\mathbf{x}-\mathbf{x}_{i}(t))^{\top}\mathbf{H}^{-1}(\mathbf{x}-\bm{x}_{i}(t))\right),$ (B.1)
where $\mathbf{H}$ is the symmetric positive-definite bandwidth matrix, $\mathbf{x}\in\mathbb{R}^{2}$ is the spatial domain and $\bm{x}_{i}(t)\in\mathbb{R}^{2}$ is the position of the $i$-th pedestrian.
@@
$h_{i}=\sigma_{i}\left(\frac{4}{N(d+2)}\right)^{1/(d+4)},$ (B.2)
@@
$\Omega^{\text{ext}}=[a-\Delta x,b+\Delta x]\times[c,d],$ (B.3)
while the support in the $y$-direction remains unbounded. To enforce continuity and reduce boundary-induced artifacts in the $x$-direction, we further augment the observed positions with “ghost” particles near the domain edges. Let $\epsilon=\Delta x$ define a threshold for proximity to the boundary. For each position $\bm{x}_{i}(t)=(x_{i}(t),y_{i}(t))$ such that $x_{i}(t)\leq a+\epsilon$, we introduce a “mirrored” pedestrians at
$\bm{x}_{i}^{\text{aug1}}(t)=(x_{i}(t)-(b-a),y_{i}(t)).$
Similarly, for every $\bm{x}_{i}(t)$ such that $x_{i}(t)\geq b-\epsilon$, we introduce a “mirrored” pedestrians at
$\bm{x}_{i}^{\text{aug2}}(t)=(x_{i}(t)+(b-a),y_{i}(t)).$
@@
$\{\bm{x}_{i}^{\prime}(t)\}=\{\bm{x}_{i}(t)\}\cup\{\bm{x}_{i}^{\text{aug1}}(t)\}\cup\{\bm{x}_{i}^{\text{aug2}}(t)\}.$ (B.4)
A pseudo-code for the estimation of the density in $\Omega\subset\mathbb{R}^{2}$ is given in Algorithm B.1.
Algorithm B.1 Kernel Density Estimation (KDE) for extracting density fields from pedestrian positions.
@@
4:Define extended support $\Omega^{\text{ext}}$ to handle periodic boundary conditions $\triangleright$ Use Eq. (B.3)
@@
6:Introduce “mirrored” pedestrians to form $\{\mathbf{x}_{i}^{\prime}(t)\}$ $\triangleright$ Use Eq. (B.4)
@@
9:  Compute kernel contribution $K_{h}(\mathbf{x}-\mathbf{x}_{p}(t))$ over the extended support $\Omega^{\text{ext}}$ $\triangleright$ Use Eq. (B.1)
@@
14:Employ binary mask function to ensure zero density in obstacle subdomain $\Gamma$, by setting $\triangleright$ See Eq. (B.5)
@@
$M(x,y)=\begin{cases}0,&\text{if }(x,y)\in\Gamma\subset\Omega,\\ 1,&\text{otherwise}.\end{cases}$ (B.5)